OpenStudy (anonymous):

I will give medal!!! Write a two-step equation that needs the steps of addition then multiplication in order to be solved, and then solve it. Show your work.

OpenStudy (one098):

Do you know what a two-step equation is?

OpenStudy (anonymous):

4x-6=2

OpenStudy (anonymous):

4x-6=2 +6 +6

OpenStudy (anonymous):

4x=8

OpenStudy (anonymous):

4x=8 /4 /4

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

or if its multiplication 4(x-6)=2 4x-24=2 +24=+24 4x=26 x=26/4 x=6.5

OpenStudy (anonymous):

but the best possible answer would be x-6/4=2

OpenStudy (anonymous):

4(x-6/4)=(2)4

OpenStudy (anonymous):

x-6=4

OpenStudy (anonymous):

x=10

OpenStudy (anonymous):

thanks @FLAMER123

OpenStudy (anonymous):

No problem.